# On the Theory of Risk-Aware Agents: Bridging Actor-Critic and Economics

### v2

@@ Appendix E Experimental Details
E.1 Low and High Replay Regimes
We consider two replay regimes:
  • Compute-Efficient Regime - each algorithm performs 2 gradient update loops per environment step.
  • Sample-Efficient Regime - each algorithm performs 16 gradient update loops per environment step, with full-parameter resets every $160k$ environment steps. This configuration is based on the one described in (D’Oro et al., 2022).
The evaluation spans $1$mln environment steps on $21$ tasks listed in Table 2. The set of tested algorithms includes:
@@ E.1 Low and High Replay Regimes
  • SAC - Soft Actor-Critic (Haarnoja et al., 2018) builds on DDPG (Silver et al., 2014). SAC extends the DDPG algorithm with features such as a stochastic policy using the reparametrization trick, automatic entropy temperature adjustment, Clipped Double Q-learning, and maximum entropy updates for actor and critic networks.
@@
  • TOP - Tactical Optimism and Pessimism (Moskovitz et al., 2021) addresses the optimism-pessimism dilemma with an external discrete bandit learning $\beta^{p}$ from predefined values. Initially based on TD3, the implementation here uses SAC due to its better performance.
  • GPL - Generalized Pessimism Learning (Cetin & Celiktutan, 2023) proposes updating pessimism in the pessimistic value approximation $V^{\pi}_{\beta}$ through a dual optimization objective.
We standardize the common hyperparameters between all algorithms to minimize the differences between all implementations. We align the basic values with the state-of-the-art SAC implementation (D’Oro et al., 2022). We employ uniform network architectures and a standard ensemble of two critics (Fujimoto et al., 2018; Haarnoja et al., 2018; Ciosek et al., 2019; Moskovitz et al., 2021; Cetin & Celiktutan, 2023). We report the detailed hyperparameter values in Appendix F. The results for these experiments are presented in Figures 1, 4, 6 and 14.
Table 2: $21$ tasks used for the evaluation of model-free algorithms in low and high replay ration regimes.
Low and High Replay Ratio
DeepMind Control MetaWorld
Acrobot (Swingup) Assembly
Dog (Run) Box (Close)
Fish (Swim) Coffee (Pull)
Hopper (Hop, Stand) Drawer (Open)
Humanoid (Run, Stand, Walk) Hammer
Pendulum (Swingup) Lever (Pull)
Swimmer (Swimmer6) Push
Walker (Run) Stick (Pull, Push)
Sweep
E.2 Comparison Against Model-Based
We evaluate the following set of continuous-action model-based and model-free algorithms:
  • DreamerV3 - a state-of-the-art model-based algorithm (Hafner et al., 2023). It was shown to perform well across a wide range of tasks, including DMC.
@@ E.2 Comparison Against Model-Based
  • D4PG - Distributed Distributional Deterministic Policy Gradients (Barth-Maron et al., 2018) is a distributed version of DDPG leveraging distributional critics, prioritized experience buffer and N-step returns.
  • MPO - Maximum a Posteriori Policy Optimisation (Abdolmaleki et al., 2018) is an off-policy algorithm based on coordinate ascent on a relative entropy objective.
For those algorithms, we source the original results reported by the authors of the respective method. In this setup, we consider two sets of environments:
  • Easy to medium tasks - this set includes tasks chosen based on a common intersection of all results reported at $500k$ environment steps by the algorithms under consideration.
  • Complex tasks - This set comprises $5$ of the most challenging tasks from the DMC suite, known to be solvable by pure RL algorithms. For these complex tasks, the comparison is specifically against TD-MPC, as it was shown to reliably solve dog tasks with proprioceptive input, as per (Hansen et al., 2022). To better evaluate performance, a prolonged training duration of $3$mln environment steps is implemented, with an action repeat of $2$, consistent with the original TD-MPC setup. Additionally, to minimize the architecture complexity gap between DAC and TD-MPC, DAC employs the TD-MPC critic architecture, including an extra hidden layer and the use of layer normalization.
All tasks included in these experiments are detailed in Table 3. The results of these experiments are showcased in Figures 5, 16 and 16.
Table 3: $15$ tasks used for the evaluation of DAC against model-based algorithms.
Comparison Against Model-Based
$10$ Easy to Medium (500k steps) $5$ Complex (3mln steps)
Acrobot (Swingup) Dog (Run, Trot, Walk)
Cartpole (Swingup-Sparse) Humanoid (Run, Walk)
Cheetah (Run)
Finger (Spin, Turn-Hard)
Hopper (Hop)
Reacher (Easy, Hard)
Walker (Run, Walk)
E.3 Hyperparameter Sensitivity & Design Choices
In the experiments addressing hyperparameter sensitivity and design choices, we evaluate the configurations on a selected set of $15$ tasks, as outlined in Table 4. These tasks represent a subset of the $21$ tasks used in the main experiment. Each configuration is executed for a duration of $500k$ environment steps.
E.3.1 Hyperparameter Sensitivity
In our study on the hyperparameter sensitivity of the proposed approach, we focus on several key hyperparameters, examining their impact on the performance of the algorithm. We list the introduced hyperparameters along with the tested values below, with the results for this evaluation is presented in Figure 7.
  • KL divergence target ($\mathcal{KL}^{*}$) - this parameter represents the desired KL divergence between the pessimistic and optimistic policies. When the observed KL divergence exceeds this target, DAC adjusts by decreasing $\beta^{o}$ (reducing the risk appetite of the optimistic actor) and increasing $\tau$ (enhancing the influence of KL regularization). The KL target is defined per action dimension, as detailed in Equation 20. The values tested for $\mathcal{KL}^{*}$ are: $0.05$, $0.25$, $0.5$, and $0.75$.
  • Standard deviation multiplier ($m$) - this hyperparameter determines the ratio between the standard deviations of the optimistic and pessimistic policies. For instance, if $m=2$, the standard deviation targeted by the optimistic policy will be double that of the pessimistic policy, as defined in Equation 44. The considered values for $m$ are: $0.75$, $1.0$, $1.25$, and $1.5$.
  • Adjustment mechanism optimizer settings - the adjustment mechanism for $\beta^{o}$ and $\tau$ is gradient-based and thus requires specific optimizer settings, including the learning rate. The same optimizer settings are applied to both $\beta^{o}$ and $\tau$. The learning rates evaluated are: $53e-$ 5, $43e-$ 4, $33e-$ 3, and $23e-$ 2.
E.3.2 Design Choices
In this section, we explore various simplifications of DAC design to assess their impact on performance. These modifications involve alterations to the structure and behavior of the actors in the DAC framework. The specific design simplifications considered are:
  • Single policy $\pi_{\eta}^{o}$ this variant employs SAC with only an optimistic actor, thus relying solely on the optimistic policy for exploration, TD learning, and evaluation. Due to the presence of a single actor, the optimism adjustment mechanism is not applicable in this setup.
  • Deterministic $\pi_{\phi}^{p}$ DAC is modified to have a completely deterministic pessimistic actor, while the optimistic actor remains stochastic and responsible for exploration. This setup often leads to an overestimation of Q-values, indicating the potential regularizing effect of soft SARSA.
  • Deterministic $\pi_{\eta}^{o}$ in this configuration, the optimistic actor of DAC is made fully deterministic. Consequently, the exploration policy becomes deterministic and greedily maximizes the Q-value upper bound.
  • DAC simplified - this refers to the streamlined version of DAC, as described in Section 4.
  • No KL regularization - this DAC variant removes the KL penalty on the optimistic actor, allowing it to diverge from the conservative actor without constraints.
  • No $\tau$ and $\beta^{o}$ adjustment - this variant operates without automatic adjustments to the rate of optimism and the KL weight, keeping these parameters fixed throughout training.
  • No $\beta^{o}$ adjustment - DAC is modified to retain the automatic adjustment of the KL weight, but without the automatic adjustment of optimism.
  • No $\tau$ adjustment - DAC without the automatic adjustment of the KL weight, albeit still using the automatic adjustment of the optimism.
The findings from the DAC design simplifications, as detailed in Table 1, reveal some intriguing insights into the dynamics of actor behavior within the DAC framework. Firstly, the variant with a deterministic pessimistic actor exhibits relatively poor performance. Interestingly, it underperforms compared to the configuration with a deterministic optimistic actor, which employs deterministic greedy exploration. This outcome suggests that incorporating variance in TD learning may offer benefits that warrant further investigation. The deterministic pessimistic actor’s subpar performance indicates the potential significance of stochastic elements in policy learning. Furthermore, in low replay settings, the simplified DAC design performs comparably to the regular DAC model. However, a notable performance gap emerges in high replay scenarios. This finding indicates that the regular architecture may be more beneficial when the learning is more aggressive. Finally, the results demonstrate a synergistic effect when both $\beta^{o}$ (optimism rate) and $\tau$ (KL weight) are subject to adjustment mechanisms. This synergy leads to performance improvements that surpass those achieved by adjusting either parameter individually. This suggests that the interplay between optimism in policy and KL regularization is crucial for balancing exploration and exploitation, and for achieving optimal performance. These insights highlight the importance of carefully considering the balance between deterministic and stochastic elements in actor designs.
Table 4: $15$ tasks used for the evaluation of DAC hyperparameter sensitivity and design choice evaluation.
Hyperparameter Sensitivity & Design Choices
DeepMind Control MetaWorld
Acrobot (Swingup) Box (Close)
Fish (Swim) Drawer (Open)
Hopper (Hop, Stand) Hammer
Humanoid (Run, Stand, Walk) Push
Pendulum (Swingup) Sweep
Swimmer (Swimmer6)
Walker (Run)
Appendix F Learning Curves and Hyperparameters
@@
Table 5: We take hyperparameter values proposed by the authors of the respective method.
@@
DAC Pessimism $-0.33$
Initial Optimism $0.5$
@@
Figure 14: Learning curves for the compute-efficient ($2$ gradient steps per environment step) and sample-efficient ($16$ gradient steps per environment step and full-parameter resets every $160k$ environment steps) benchmarks. $Y$-axis reports mean with $3$ standard deviations. $10$ seeds per task.
Figure 15: Learning curves for the easy to medium tasks taken from the DreamerV3 and TD-MPC evaluations. $X$-axis denotes the environment steps in thousands; and $Y$-axis reports mean with $3$ standard deviations. $10$ seeds per task.
Figure 15: Learning curves for the easy to medium tasks taken from the DreamerV3 and TD-MPC evaluations. $X$-axis denotes the environment steps in thousands; and $Y$-axis reports mean with $3$ standard deviations. $10$ seeds per task.
Figure 16: Learning curves for the complex tasks taken TD-MPC evaluations. To assure the consistency of results we use action repeat of $2$ (ie. the value used by TD-MPC). $X$-axis denotes the environment steps in thousands; and $Y$-axis reports mean with $3$ standard deviations. $5$ seeds per task.

### v3

@@ Appendix E Experimental Details
E.1 Environments
The table below lists task used in the main experiment.
Table 2: $30$ tasks used for the evaluation of model-free algorithms in low and high replay ration regimes. The $15$ bold tasks are used for the hyperparameter sensitivity and design choices ablation.
DeepMind Control MetaWorld MyoSuite
Acrobot Swingup Assembly Reach Easy
Cheetah Run Box Close Reach Hard
Hopper Hop Coffee Pull Pose Easy
Humanoid Run Drawer Open Pose Hard
Humanoid Stand Hammer Pen Twirl Easy
Humanoid Walk Lever Pull Pen Twirl Hard
Swimmer Swimmer6 Push Object Hold Easy
Pendulum Swingup Stick Pull Object Hold Hard
Quadruped Run Stick Push Key Turn Easy
Walker Run Sweep Key Turn Hard
E.2 Baseline Algorithms
We standardize the common hyperparameters between all algorithms to minimize the differences between all implementations. We align the basic values with the state-of-the-art SAC implementation (D’Oro et al., 2022), thus limiting the requirement for extensive hyperparameter search across algorithms. We employ uniform network architectures and a standard ensemble of two critics (Fujimoto et al., 2018; Haarnoja et al., 2018a; Ciosek et al., 2019; Moskovitz et al., 2021; Cetin & Celiktutan, 2023). The set of tested algorithms includes:
@@ E.2 Baseline Algorithms
  • SAC - Soft Actor-Critic (Haarnoja et al., 2018a) builds on DDPG (Silver et al., 2014). SAC extends the DDPG algorithm with features such as a stochastic policy using the reparametrization trick, automatic entropy temperature adjustment, Clipped Double Q-learning, and maximum entropy updates for actor and critic networks.
@@
  • TOP - Tactical Optimism and Pessimism (Moskovitz et al., 2021) addresses the optimism-pessimism dilemma with an external discrete bandit learning $\beta^{p}$ from predefined values. Based on TD3.
  • GPL - Generalized Pessimism Learning (Cetin & Celiktutan, 2023) proposes updating pessimism in the pessimistic value approximation $V^{\pi}_{\beta}$ through a dual optimization objective. Based on SAC.
@@
In high replay ratio, all algorithms undergo full-parameter resets according to the schedule proposed by D’Oro et al. (2022). As such, given replay ratio of 16 gradient steps per environment step, the agents are resetted in $(160k,320k,480k,640k,800k)$ steps. ma
@@
Table 3: For non-shared hyperparameters we take values which were found to perform best by the authors of the respective method.
@@
DAC Pessimism $-0.2$
Initial Optimism $1.0$
@@
E.3 Training Curves
Figure 15: Training curves for DMC ($RR=2$ rows $1$ $2$, $RR=16$ row $3$ $4$). $Y$-axis denotes IQM and $X$-axis denotes environment steps. 10 seeds per task.
Figure 16: Training curves for MW ($RR=2$ rows $1$ $2$, $RR=16$ rows $3$ $4$). $Y$-axis denotes IQM success rates and $X$-axis denotes environment steps. 10 seeds per task.
Figure 17: Training curves for MYO ($RR=2$ rows $1$ $2$, $RR=16$ rows $3$ $4$). $Y$-axis denotes IQM success rates and $X$-axis denotes environment steps. 10 seeds per task.